OpenStudy (anonymous):

Can anyone help me with (sec(x)+tan(x)/sec(x)-tan(x))=(sec(x)+tan(x))^2

OpenStudy (aravindg):

Open up (sec(x)+tan(x))^2

OpenStudy (aravindg):

Then use the identity sec^2 x+tan^2 x=1

OpenStudy (anonymous):

alright so i now have the left side = to 1

OpenStudy (anonymous):

do i multiply by secx+tanx top and bottom on the left?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sec^2 x-tan^2 x=1 and not +

OpenStudy (anonymous):

i then have sec(x)^2+2sec(x)tan(x)+ tan(x)^2 / sec^2-tan^2

OpenStudy (anonymous):

It is straight forward now \[ \frac{(\tan (x)+\sec (x)) (\tan (x)+\sec (x))}{(\sec (x)-\tan (x)) (\tan (x)+\sec (x))}=\frac{(\tan (x)+\sec (x))^2}{\sec ^2(x)-\tan ^2(x)}=\ (\tan (x)+\sec (x))^2 \]

OpenStudy (anonymous):

\[ \frac{(\tan (x)+\sec (x)) (\tan (x)+\sec (x))}{(\sec (x)-\tan (x)) (\tan (x)+\sec (x))}=\frac{(\tan (x)+\sec (x))^2}{\sec ^2(x)-\tan ^2(x)}=\\ (\tan (x)+\sec (x))^2 \]

OpenStudy (anonymous):

thank you i had foiled the problem and gotten confused

OpenStudy (anonymous):

It is pretty simple